OpenStudy (anonymous):

(5/(x^3+5x^2))=(4/(x+5))+(1/x^2) I found the extraneous solution for 0, and I need to solve for x. I want to isolate x out of the denominator, what can I multiply this with?

OpenStudy (anonymous):

I don't know what I can multiply all of these with so that the denominator disappears

myininaya (myininaya):

first of all, can x^3+5x^2 be factored?

OpenStudy (anonymous):

yes into x^2(x+5)

myininaya (myininaya):

\[\frac{5}{x^2(x+5)}=\frac{4}{x+5}+\frac{1}{x^2}\] Second of all, we definitely no x cannot be what two numbers? --- third of all, what is the least common multiply of (x+5) and x^2? --- fourth of all, the lcm is what we will choose to multiply both sides by.

myininaya (myininaya):

So second question, we know x cannot be what two numbers?

OpenStudy (anonymous):

0 and -5 right?

myininaya (myininaya):

Right on! :)

myininaya (myininaya):

Ok so if we get a solution when we solve this that is one of those numbers we will know that is a contradiction and it definitely is not a solution

myininaya (myininaya):

Anyways, the third question: what is the least common multiply (lcm) of x^2 and (x+5)

myininaya (myininaya):

if you don't know, sometimes it is just best to find a common multiple and you can do that pretty quick by multiplying the x^2 and the (x+5) notice these were factors in the bottom in the fractions

OpenStudy (anonymous):

the x^2 and x+5 both are x I think?

myininaya (myininaya):

lcm of x^2 and (x+5) is just x^2(x+5)

myininaya (myininaya):

and that is what we will multiply both sides by

OpenStudy (anonymous):

oh ok so do I multiply that out of all the denominators?

myininaya (myininaya):

\[\frac{x^2(x+5)}{1} \cdot \frac{5}{x^2(x+5)}=\frac{x^2(x+5)}{1} \cdot \frac{4}{x+5}+\frac{x^2(x+5)}{1} \cdot \frac{1}{x^2}\] cancel away

OpenStudy (anonymous):

i got 5=4x^2+x^2

myininaya (myininaya):

the only one I'm worried about is your last term there

OpenStudy (anonymous):

sorry I replied before looking at your image

myininaya (myininaya):

its k it looks almost good just the last term needs to be corrected

OpenStudy (anonymous):

sorry last one is x+5

OpenStudy (anonymous):

messed that up

myininaya (myininaya):

yep yep :)

myininaya (myininaya):

so you need to solve 5=4x^2+x+5

OpenStudy (anonymous):

Thanks ! I got 4x^2+x=0 sorry I'm kinda stuck right here.... how should I simplify this? I know it's a ax^2+bx+c form...

myininaya (myininaya):

4x^2+x=0 is not too scary you can actually factor the 4x^2+x

myininaya (myininaya):

both 4x^2 and x have what as a common factor?

OpenStudy (anonymous):

x

myininaya (myininaya):

so 4x^2+x=x times (what? )

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

4x(x)=4x^2 right?

myininaya (myininaya):

\[4x^2+x=x(\frac{4x^2}{x}+\frac{x}{x})=x(4x+1)\]

OpenStudy (anonymous):

how come there is a +1 at the end?

myininaya (myininaya):

factor an x from each term you need to divide each term by x

OpenStudy (anonymous):

wait I figured it out sorry.

myininaya (myininaya):

x times (4x+1)=4x^2+x

myininaya (myininaya):

ok

myininaya (myininaya):

so you know then we have 4x^2+x=0 which we rewrote as x(4x+1)=0

myininaya (myininaya):

x(4x+1)=0 means we have either x=0 or 4x+1=0 or both=0

myininaya (myininaya):

so what is our solution

OpenStudy (anonymous):

so I might need a little more help on this: I know in order to find the solution I need the x isolated.

myininaya (myininaya):

ya

myininaya (myininaya):

x=0 (we already know this isn't a solution because remember we said x could not be 0 at the beginning) --- so we only need to solve 4x+1=0

OpenStudy (anonymous):

oh i see then it's x=-1/4

myininaya (myininaya):

yes

myininaya (myininaya):

that is correct good job nick

myininaya (myininaya):

you were able to answer a lot of questions on this problem that I gave you did really well

OpenStudy (anonymous):

thank you... so to sum it up, the extraneous solutions are only 0, -5 and the x is -1/4? Thank you so so so so so much!

myininaya (myininaya):

well the extraneous solutions was only x=0 (I did not include x=-5 in the extraneous solution set because we did not get -5 in the end) but the only solutions (non-extraneous) is x=-1/4

OpenStudy (anonymous):

but technically -5 is an extraneous solution, correct?

myininaya (myininaya):

http://www.mathwords.com/e/extraneous_solution.htm "A solution of a simplified version of an equation that does not satisfy the original equation.-extraneous solution " This is the definition I'm using. 4x^2+x=0 was a simplified version of the original equation (the equation we started with) this simplified equation 4x^2+x=0 didn't give us x=-5 so I do not include in the extraneous solution set I only include the x=0 in the extraneous solution set but the actual solution to the problem is x=-1/4

OpenStudy (anonymous):

Oh I see what you mean thanks so much again! I wouldn't have been able to do this without you! :P

myininaya (myininaya):

Np. http://hotmath.com/hotmath_help/topics/extraneous-solutions.html Here is another example. Notice in that first problem there they do not call x=-2 an extraneous solution but they only call x=2 an extraneous solution because it works for the simplified version but not the original version.

OpenStudy (anonymous):

Ill take a look at it. Thanks!

myininaya (myininaya):

I will give you a medal for your awesome work into this problem. You did super.